OpenStudy (anonymous):

I'm having a problem figuring out the difference in terms of an anti-derivative and integral. The problem is:

OpenStudy (anonymous):

Anti-derivatives are used to solve integrals.

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ x dx }{ (7x^2+3)^5 }\] I'm given this problem on two equations. One asks for me to evaluate the integral and the other asks me to find the anti-derivative. Using the sub method, I get \[-\frac{ 1 }{ 56 }(7x^2+3)^(-4)+ C\] for both problems. Can someone explain to me what I'm doing wrong

OpenStudy (anonymous):

Uhh. That should be negative 1/56 times( 7x to the 2nd power + three in parentheses to the negative 4th power) + C

OpenStudy (anonymous):

Evaluating an improper integral is just finding the anti-derivative.

Miracrown (miracrown):

Your result here is correct. Note that integral and anti-derivative are synonyms.

Miracrown (miracrown):

You then correctly made use of a U substitution to do the integral.

OpenStudy (anonymous):

Thanks! I'll keep in mind that the two are synonyms. That really helped. Appreciate the help from both of you!

Miracrown (miracrown):

yw :)

OpenStudy (anonymous):

They aren't synonymous, they are just closely related.